OpenStudy (anonymous):

what is the highest common multiple of 16 and 24?

OpenStudy (anonymous):

Split each number up into its prime factors: \[16=2\times2\times2\times2=2^4\] \[24=2\times2\times2\times3=2^3\times3\] Both have \(2^3\) as a factor, so the highest common multiple is \(2^3=8\).

OpenStudy (anonymous):

Do you understand that or is there anything I can make clearer?

OpenStudy (anonymous):

thanks for your answer!

OpenStudy (anonymous):

No problem :)

OpenStudy (amistre64):

ive heard of a lowest common multiple; and a highest common factor ...

OpenStudy (amistre64):

but there is no highest common multiple of a set of integers that i can recall

OpenStudy (anonymous):

I think he meant HCF.

OpenStudy (amistre64):

with any luck :)

OpenStudy (anonymous):

Highest common multiple would be \(\infty\) I guess :P

OpenStudy (anonymous):

my teacher says the highest common multiple, but they also called it HCF, but i just called it highest common factor because thats what it says in my homework.

OpenStudy (anonymous):

Highest common factor is fine, your teacher is wrong to say highest common multiple though.I think she's confusing it with lowest common multiple.

OpenStudy (anonymous):

sorry, multiple. just got confused

OpenStudy (anonymous):

He/she :S